Shift summary for day, 2 Sept, 1999

Angela Bellavance (bellavan@rice.edu)
Thu, 2 Sep 1999 15:36:38 -0500 (CDT)

Shift Summary for day shift, 02/Sep/99

On Shift:A. Tripathi, A. Bellavance

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: No beam - shutdown for access

========================================================================
Runs ended this shift: None that were saved
Tapes brought to FCC this shift: None

========================================================================
Accesses during shift:

Jason - plugged in 2X7, upstream cable at the repeater
Sasha L. & Valeri - TRD work
Doug - DC laser positions and check pump oil.

========================================================================
Overview of shift:

Sit around and label tapes (and take notes of what else happened).

========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
7:00 Take over shift with no beam. Message on Ch.13 is "beginning
access 0700-1600".
7:31 Begin cosmic run with no writing to disk or tape.
7:45 Ed calls to check up on us. Not much to check up on.
8:42 Alarm sounds: KTeV detector temp is 0.0F... Ignored.
9:26 Same alarm sounds again, ignored again.
? Angela and BobT are shocked to learn that BobH has never seen "The
Princess Bride".
11:40 Access by JasonL, SashaL, and Valeri starts.
11:55 Jason is out of access.
? Sasha and Valeri out of access.
12:50 Crew from maintenance makes access.

Other events:

JasonL has put the X-view fruit back into the readout and the plan is to
leave it in.

Somebody stopped by to look at the Bry air supply. It was not turned off.

Crew from maintenance looks at exhaust fans downstairs.

DougJ went on access as requested by Ed to do a position scan of the
chambers and check the pump oil (was fine, no fill).

RickK requests we initialize a run called BEAM_832_KE3HOG2. During
initialization the trigger went into error state, but don't know if it's
related. Otherwise it initialized fine.

Shift crew (with help from JasonH) labels a few E832 tapes and 220 E799
tapes.

========================================================================
Follow up Notes on Previous Shift:

I'm not sure if the previous shift sent out a shift summary. We didn't
find anything lurking in an editor anywhere like before so there's not
much we can do about it.

========================================================================
Items to Follow up on Next Shift

The magnet is still off and will need to be turned back on.

Most of the other enclosure magnets have tripped off during the day. They
should all be checked and turned back on once accesses are done.

There are currently no tapes in the tape drives.

Sometime in the future, a KTeV movie night should be instituted to correct
the theatrical ignorance of KTeV members. :)

For Arun, Angela

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke